OpenStudy (antonialynn):

4 - 3 ln x = -8 I'm really not getting this math thing , y'all

Nnesha (nnesha):

\[\large\rm 4-3 \color{REd}{lnx} = -8\] can you solve for `lnx` ??:=))

OpenStudy (antonialynn):

is the answer 4?

OpenStudy (antonialynn):

thats wha† i got

Nnesha (nnesha):

how did you get 4?

Nnesha (nnesha):

ahh btw we have to solve for x therfore we have to take `e` both sides in order to cancel out ln but first get rid of all the numbers that are on the left side \[e^{\ln x} = x\]

Nnesha (nnesha):

jst solving for `lnx` wouldn't give u the final answer.

OpenStudy (antonialynn):

im so confused :(

OpenStudy (kingnoob):

You're on the right track so far, just need to finish it.. I'm assuming you ended up with \(\large lnx =4\) ?

OpenStudy (antonialynn):

Yessss

OpenStudy (kingnoob):

Now use what Nnesha suggested earlier with "e" (:

OpenStudy (antonialynn):

but i dont get where the e comes in? or what im doing with it

Nnesha (nnesha):

okay thats fine \[\large\rm 4-3lnx =-8\] first subtract 4 both sides and then divide by -3  \[\frac{ - 3\ln x }{ -3}=\frac{-12}{-3} \rightarrow lnx=4\] now rewrite both sides as an exponnet of e